OpenStudy (anonymous):

4x-2z=3y+8 (if y=12 and z = -3) solve for x

OpenStudy (anonymous):

just plug the values of y and z \[4x-2(-3)=3(12)+8\] \[4x+6=36+8\] now solve for x

OpenStudy (anonymous):

thanks!